# Extendability of projective varieties via degeneration to ribbons with applications to Calabi-Yau threefolds

### v1

@@ Abstract.
In this article we study the extendability of a smooth projective variety by degenerating it to a ribbon. We apply the techniques to study extendability of Calabi-Yau threefolds $X_{t}$ that are general deformations of Calabi-Yau double covers of Fano threefolds of Picard rank $1$. The Calabi-Yau threefolds $X_{t}\xhookrightarrow{|lA_{t}|}\mathbb{P}^{N_{l}}$, with $l\geq j$, where $A_{t}$ is the generator of Pic$(X_{t})$ and $j$ is the index of $Y$, are general elements of a unique irreducible component $\mathcal{H}_{l}^{Y}$ of the Hilbert scheme which contains embedded Calabi-Yau ribbons on $Y$ as a special locus. For $l=j$, using the classification of Mukai varieties, we show that the general Calabi-Yau threefold parameterized by $\mathcal{H}_{j}^{Y}$ is as many times smoothly extendable as $Y$ itself. On the other hand, we find for each deformation type $Y$, an effective integer $l_{Y}$ such that for $l\geq l_{Y}$, the general Calabi-Yau threefold parameterized by $\mathcal{H}_{l}^{Y}$ is not extendable. These results provide a contrast and a parallel with the lower dimensional analogues; namely, $K3$ surfaces and canonical curves ([CLM93], [CLM93A], [CLM98]) which stems from the following result we prove: for $l\geq l_{Y}$, the general hyperplane sections of elements of $\mathcal{H}_{l}^{Y}$ fill out an entire irreducible component $\mathcal{S}_{l}^{Y}$ of the Hilbert scheme of canonical surfaces which are precisely $1-$ extendable with $\mathcal{H}^{Y}_{l}$ being the unique component dominating $\mathcal{S}_{l}^{Y}$. The contrast lies in the fact that for polarized $K3$ surfaces of large degree, the canonical curve sections do not fill out an entire component while the parallel is in the fact that the canonical curve sections are exactly one-extendable.
@@ 1. Introduction
The extendability of a projective variety is a natural and fundamental question in projective geometry. This classical question has been the topic of intense research for decades and has revealed deep connections between the geometry of the embedding, Gaussian map of curve sections, deformations of cones over the hyperplane sections etc. We refer to [Lop23] for an excellent recent survey of this topic. The purpose of this article is to introduce techniques to study extendability of smooth projective varieties via degeneration to ribbons. In particular we show two different themes, one to show extendability and the other to show non-extendability of the general member of a one parameter family that projectively degenerates to a ribbon in the central fiber. We apply them to study extendability of Calabi-Yau threefolds, where not much is known. We will now recall the definition of a ribbon
@@ 1.1. Extendability of the general member of a component of the Hilbert scheme containing ribbons
    Show that the extended ribbon $\widetilde{W}\subset\mathbb{P}^{M+k}$ is smoothable. Thus we have shown the $k-$ step extendability of some smooth special members of $\mathcal{H}$ containing the ribbons in their closure.
@@ 1.2. Non-extendability of the general member of a component of the Hilbert scheme containing ribbons
To show non-extendability, we apply Zak-Lv́ovsky’s criterion on the ribbon $\widetilde{Y}\subset\mathbb{P}^{N}$. We recall the definition of the invariant $\alpha(X)$ and two results by Zak and Lv́ovsky ([Lvo92], [Zak91], see also [Bad94], [Lop23, Theorem $1.3$]).
@@
([Lvo92], [Zak91]) Let $X\subset\mathbb{P}^{N}$ be a smooth irreducible nondegenerate variety of codimension at least $1$ and suppose $X$ is not a quadric. If $\alpha(X)\leq 0$, then X is not extendable. Further, given an integer $k\geq 2$, suppose that either:
@@
By results in [BE95], [Fon93], [Gon06], [GGP10], [GGP13], it is known that under sufficiently general conditions, as an embedding degenerates to a double cover, the embedded models degenerate to a ribbon over the base of the double cover. Since the ribbon is a local complete intersection, by a series of short exact sequences, it is shown that computing the cohomology of the twists of the normal bundle $N_{\widetilde{Y}/\mathbb{P}^{N}}$ of a embedded ribbon $\widetilde{Y}$ on $Y\subset\mathbb{P}^{N}$ eventually boils down to computing the cohomology of the twists of the normal bundle $N_{Y/\mathbb{P}^{N}}$ (see equations 2.1, 2.2, 2.3) which are easier to handle than the normal bundle of a general member $X\subset\mathbb{P}^{N}$ in the family, about which we know very little.
In this paper we study the extendability of Calabi-Yau threefolds $X\subset\mathbb{P}^{N}$ by degenerating it to Calabi-Yau ribbons on Fano-threefolds $Y$ of Picard rank one, embedded by the complete linear series of $-lH$ where $H$ is the generator of Pic$(Y)$. For $l$ equal to the index of $Y$, we show that the Calabi-Yau threefolds are at least as many times smoothly extendable as the Fano threefold $Y$. For higher values of $l$, we show non-extendability of $X$. Fano threefolds of Picard rank $1$ are classified into $17$ different deformation types by the classification of Iskovskih-Mori-Mukai ([Isk77], [Isk78], [MM82]). We use the explicit description of each of the families which are either complete intersections in weighted projective spaces or regular sections of homogeneous vector bundles on Grassmannians (see [DFT22], [Bel24], https://www.fanography.info/.). The computations boil down to Borel-Bott-Weil theorems. We now state our results. First we state a Proposition describing the Calabi-Yau threefolds we work with and their Hilbert schemes.
@@
    There exists a unique irreducible component $\mathcal{H}^{Y}_{l}$ of the Hilbert scheme of $\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ with Hilbert polynomial $p(l)=h^{0}(lH)+h^{0}((l-j)H)$ that parameterizes linearly normal Calabi-Yau varieties $X_{t}\subset\mathbb{P}^{N_{l}}$ such that $(X_{t},\mathcal{O}_{X_{t}}(1))$ can be deformed to $(X,lA)$ along an irreducible one parameter family.
@@
It is interesting to compare the results of this article with the results on extendability of $K3$ surfaces and canonical curves in [CLM93], [CLM93A] and [CLM98]. Let $\mathcal{H}_{r,g}$ denote the Hilbert scheme of $K3$ surfaces embedded by the $r-$ th veronese map of a primitive very ample line bundle of sectional genus $g$. By studying Gaussian maps of canonical curve sections of such $K3$ surfaces, it was proven that for $(r=1,g\geq 13),(r=2,g\geq 7),(r=3,g\geq 5),(r=4,g\geq 4),(r\geq 5,g\geq 3)$ (see [CLM98, Table $2.14$] for the precise and slightly stronger results), the general curve section of the general $K3$ surfaces in $\mathcal{H}_{r,g}$ has corank one Gaussian map which in particular means that the general $K3$ surfaces are not extendable and the general curve sections are exactly one extendable. The results Theorem 1.6, $(1)$, $(3)$ are analogues of the above results. However, there is a contrast for Calabi-Yau threefolds; the hyperplane sections of $\mathcal{H}^{Y}_{l}$ for $l\geq l_{Y}$ fill out an entire irreducible component of the Hilbert scheme of canonical surfaces, while curves on a $K3$ surface are non-general curves of the corresponding component of the Hilbert scheme. We plot the invariants of the surface sections in subsection 4.3.
@@ 2. Main result on extendability of projective varieties
    Ribbons $\widetilde{Y}\subset\mathbb{P}^{N}$ on $Y$ with conormal bundle $L$ extending a given (possibly degenerate) embedding $Y\subset\mathbb{P}^{N}$ are parameterized by the nowhere vanishing sections of $H^{0}(N_{Y/\mathbb{P}^{N}}\otimes L)$ (see [Gon06, Proposition $2.1$]).
@@
Proof. We have the following exact sequences (see [GGP08, Lemma $4.1$])
@@
Since $\alpha(X)=0$, we have that $h^{0}(N_{X/\mathbb{P}^{N+1}}(-2))=0$ (see [Lop23, Proof of Theorem $1.3$, Claim $2.3$]). Combined with the fact that $h^{1}(N_{X/\mathbb{P}^{N+1}}(-3))=0$, we have that $h^{0}(N_{Z/\mathbb{P}^{N}})(-2)=0$. Now $h^{0}(N_{Z/\mathbb{P}^{N}})(-1)-N-1=\alpha(Z)$. Hence the dimension of the tangent space to the Hilbert scheme at the point $C(Z)$ is given by $h^{0}(N_{Z/\mathbb{P}^{N}})+\alpha(Z)+N+1=h^{0}(N_{Z/\mathbb{P}^{N}})+\alpha(X)+N+2=h^{0}(N_{Z/\mathbb{P}^{N}})+h^{0}(N_{X/\mathbb{P}^{N+1}}(-1))$.
@@ 3. Calabi-Yau ribbons on Fano threefolds and deformations of Calabi-Yau double covers of Fano threefolds
    There exists a unique irreducible component $\mathcal{H}^{Y}_{l}$ of the Hilbert scheme of $\mathbb{P}^{N_{l}}$ with Hilbert polynomial $p(l)=h^{0}(lH)+h^{0}((l-j)H)$ that parameterizes linearly normal Calabi-Yau varieties $X_{t}\subset\mathbb{P}^{N_{l}}$ such that $(X_{t},\mathcal{O}_{X_{t}}(1)=lA_{t})$ can be deformed to $(X,lA)$ along an irreducible one parameter family.
@@
Proof. Let us first show the values of $j$. For families $1.2-1.10$, $1.12-1.17$, this is clear by [Bel24] since $-K_{Y}$ is very ample. For $1.1$, $H=-K_{Y}$. $Y$ is a double cover of $\mathbb{P}^{3}$ branched along a sextic. We have $h^{0}(-2K_{Y})=h^{0}(\mathcal{O}_{\mathbb{P}^{3}}(2)$ and hence the morphism induced by $|-2K_{Y}|$ is $2:1$ onto the second Veronese embedding of $\mathbb{P}^{3}$. Further one can check that $-3K_{Y}$ is in fact projectively normal. So $j=3$. For $1.11$, by [Bel24], $j\geq 3$. Now by [NO02, Theorem $0.2$], $lH$ is very ample on $Y$ for $l\geq 3$, which implies $j=3$.
Now we show that if $l<j$, then $lA_{t}$ is not very ample. For family $1.1$, $l=1$, if $A_{t}$ is very ample, then it embeds $X_{t}$ as a degree $4$ Calabi-Yau threefold inside $\mathbb{P}^{4}$. This gives a contradiction. For family $1.11$, $l=1$, $A_{t}$ is not very ample since $h^{0}(A_{t})=3$. For $l=2$, if $|2A_{t}|$ is very ample, then $X_{t}$ is an aCM Calabi-Yau threefold of degree $16$ inside $\mathbb{P}^{6}$. But such a Calabi-Yau does not exist due to [KK16, Theorem $3.2$]. For family $1.12-1.15$, $j=2$. For $l=1$, $A_{t}$ is not very ample. To see this first note that $h^{0}(A_{t})=h^{0}(H)$. Now for $1.12$, $h^{0}(A_{t})=4$ and the morphism given by $|A_{t}|$ is $4:1$ onto $\mathbb{P}^{3}$. For $1.13-1.14$, if $A_{t}$ is very ample, then $|A_{t}|$ embeds $X_{t}$ inside $\mathbb{P}^{5}$. By [BC83, Remark $11$], all Calabi-Yau threefolds inside $\mathbb{P}^{5}$ are complete intersections of either two cubics or a quadric and a quartic or a quintic and a hyperplane. For $1.13$, $A_{t}^{3}=6$, which cannot be the degree of a Calabi-Yau inside $\mathbb{P}^{5}$. For $1.14$, $A_{t}^{3}=8$, so $X_{t}$ has to be the complete intersection of a quadric and a quartic. But then this is not possible by comparing $h^{1,2}(X_{t})$ with $h^{1,2}$ of the complete intersection. For $1.15$, if $A_{t}$ is very ample then $|A_{t}|$ embeds $X_{t}$ as a degree $10$, aCM Calabi-Yau threefold inside $\mathbb{P}^{6}$. By [KK16, Theorem $3.2$, Section $6$], such a Calabi-Yau threefold cannot exist. For family $1.16$, $j=3$. For $l=1,2$, $lA_{t}$ is not very ample. Note first that $H^{1}(T_{Y}\otimes K_{Y})=H^{1,2}(Y)=0$. Now since $h^{0}(lA)=h^{0}(lH)$, the morphism given by $|A|$ on the double cover $X$ factors through a $2:1$ map onto $Y$ followed by $Y$ embedded by $|lH|$. This together with the fact that $H^{1}(\mathcal{O}_{X})=0$, implies by [Weh86, Corollary $1.1$] and arguments similar to [GGP16, Theorem $1.7$] the morphism by $|lA_{t}|$ factors as a double cover of $Y$ as well. For family $1.17$, $j=4$. For $l=1,2,3$, the map given by $|A_{t}|$ is not very ample by exactly same arguments as in part $(5)$. The last statement follows by combining Lemma 3.2 and Proposition 3.1. $\blacksquare$
@@
Noting that $H^{0}(lH\otimes K_{Y})=H^{0}((l-j)H)$ we have $N^{\prime}+1=h^{0}(lH)+h^{0}((l-j)H)=N_{l}+1$. We show that all such ribbons $\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ have represent smooth points of the Hilbert scheme. This can be checked using [BGM24, Lemma $2.14$] and the exact sequence
@@
Now we prove part $(1)$. Consider the Calabi-Yau double cover $\pi:X\to Y$ branched along $B\in|-2K_{Y}|$. By [Par91], Lemma $4.2$, we have $H^{1}(\Omega_{X}^{1})=H^{1}(\pi_{*}\Omega_{X}^{1})=H^{1}(\Omega_{Y})\bigoplus H^{1}(\Omega_{Y}(\log(B))\otimes K_{Y})$. Now using the exact sequence
@@
Therefore $H^{1}(\Omega_{Y}(\log(B))\otimes K_{Y})=0$ implies $\pi^{*}$ is an isomorphism, $h^{1,1}(X)=1$ and $\textrm{Pic}(X)=\mathbb{Z}$. Let $A$ denote the generator. We have that for $X_{t}$, a general deformation of $X$, $h^{1,1}(X_{t})=1$. On the other hand, by [Deb01, Corollary 4.18] and [KMM92, Theorem 0.1], $Y$ is simply connected, so, by the Lefschetz theorem for homotopy groups (see [Laz04, Theorem 3.1.21]), so is $B$, where $B$ is the ample branch divisor of the double cover $X$. Now if $R\subset X$ denotes the ramification divisor, $R\cong B$ and hence $R$ is simply connected. Since $R$ is ample, once again by the Lefschetz theorem for homotopy groups, so is $X$. Therefore $X$ is simply connected and hence so is $X_{t}$. Therefore, $\textrm{Pic}(X_{t})=\mathbb{Z}$. Again since divisibility of a line bundle remains constant along a smooth family, we have that for any polarized deformation $(X_{t},A_{t})$ of $(X,A)$, $A_{t}$ generates the Picard group of $X_{t}$.
@@
As before since $Y$ is Fano and using Nakano-Vanishing theorem, one can check that both flanking terms of the previous exact sequence vanishes. This implies that $H^{1}(N_{\varphi})=0$ and deformations of $\varphi$ are unobstructed. This implies by [GGP13, Proposition $1.4$], that there is a one parameter family $\Phi_{T}:\mathcal{X}\to\mathbb{P}^{N_{l}}_{T}$ such that $\Phi_{0}=\varphi$ while $\Phi_{t}=\varphi_{t}:\mathcal{X}_{t}=X_{t}\to\mathbb{P}^{N_{l}}$ is an embedding. Further (Im $\Phi)_{0}=\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ while for $t\neq 0$, (Im $\Phi)_{t}=X_{t}\subset\mathbb{P}^{N_{l}}$. Note that since $\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ is linearly normal, we have that $X_{t}\subset\mathbb{P}^{N_{l}}$ is also linearly normal. Further we have $(\Phi_{T}^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}_{T}}(1))_{0}=\varphi^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}}(1)=\pi^{*}(lH)=lA$, and hence $(\Phi_{T}^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}_{T}}(1))_{t}=\mathcal{O}_{X_{t}}(1)=lA_{t}$ is deformation of $A$.

### v2

@@ Abstract.
In this article we study the extendability of a smooth projective variety by degenerating it to a ribbon. We apply the techniques to study extendability of Calabi-Yau threefolds $X_{t}$ that are general deformations of Calabi-Yau double covers of Fano threefolds of Picard rank $1$. The Calabi-Yau threefolds $X_{t}\xhookrightarrow{|lA_{t}|}\mathbb{P}^{N_{l}}$, with $l\geq j$, where $A_{t}$ is the generator of Pic$(X_{t})$ and $j$ is the index of $Y$, are general elements of a unique irreducible component $\mathcal{H}_{l}^{Y}$ of the Hilbert scheme which contains embedded Calabi-Yau ribbons on $Y$ as a special locus. For $l=j$, using the classification of Mukai varieties, we show that the general Calabi-Yau threefold parameterized by $\mathcal{H}_{j}^{Y}$ is as many times smoothly extendable as $Y$ itself. On the other hand, we find for each deformation type $Y$, an effective integer $l_{Y}$ such that for $l\geq l_{Y}$, the general Calabi-Yau threefold parameterized by $\mathcal{H}_{l}^{Y}$ is not extendable. These results provide a contrast and a parallel with the lower dimensional analogues; namely, $K3$ surfaces and canonical curves ([7], [8], [9]) which stems from the following result we prove: for $l\geq l_{Y}$, the general hyperplane sections of elements of $\mathcal{H}_{l}^{Y}$ fill out an entire irreducible component $\mathcal{S}_{l}^{Y}$ of the Hilbert scheme of canonical surfaces which are precisely $1-$ extendable with $\mathcal{H}^{Y}_{l}$ being the unique component dominating $\mathcal{S}_{l}^{Y}$. The contrast lies in the fact that for polarized $K3$ surfaces of large degree, the canonical curve sections do not fill out an entire component while the parallel is in the fact that the canonical curve sections are exactly one-extendable.
@@ 1. Introduction
The extendability of a projective variety is a natural and fundamental question in projective geometry. This classical question has been the topic of intense research for decades and has revealed deep connections between the geometry of the embedding, Gaussian map of curve sections, deformations of cones over the hyperplane sections etc. We refer to [29] for an excellent recent survey of this topic. The purpose of this article is to introduce techniques to study extendability of smooth projective varieties via degeneration to ribbons. In particular we show two different themes, one to show extendability and the other to show non-extendability of the general member of a one parameter family that projectively degenerates to a ribbon in the central fiber. We apply them to study extendability of Calabi-Yau threefolds, where not much is known. It is interesting to note that ribbons have appeared in the context of extendability of canonical curves and $K3$ surfaces in [6], [41], but in a way that is completely different from this article. We will now recall the definition of a ribbon
@@ 1.1. Extendability of the general member of a component of the Hilbert scheme containing ribbons
    Show that the extended ribbon $\widetilde{W}\subset\mathbb{P}^{N+k}$ is smoothable. Thus we have shown the $k-$ step extendability of some smooth special members of $\mathcal{H}$ containing the ribbons in their closure.
@@ 1.2. Non-extendability of the general member of a component of the Hilbert scheme containing ribbons
To show non-extendability, we apply Zak-Lv́ovsky’s criterion on the ribbon $\widetilde{Y}\subset\mathbb{P}^{N}$. We recall the definition of the invariant $\alpha(X)$ and two results by Zak and Lv́ovsky ([30], [43], see also [1], [29, Theorem $1.3$]).
@@
([30], [43]) Let $X\subset\mathbb{P}^{N}$ be a smooth irreducible nondegenerate variety of codimension at least $1$ and suppose $X$ is not a quadric. If $\alpha(X)\leq 0$, then X is not extendable. Further, given an integer $k\geq 2$, suppose that either:
@@
By results in [3], [13], [21], [18], [19], it is known that under sufficiently general conditions, as an embedding degenerates to a double cover, the embedded models degenerate to a ribbon over the base of the double cover. Since the ribbon is a local complete intersection, by a series of short exact sequences, it is shown that computing the cohomology of the twists of the normal bundle $N_{\widetilde{Y}/\mathbb{P}^{N}}$ of a embedded ribbon $\widetilde{Y}$ on $Y\subset\mathbb{P}^{N}$ eventually boils down to computing the cohomology of the twists of the normal bundle $N_{Y/\mathbb{P}^{N}}$ (see equations 2.1, 2.2, 2.3) which are easier to handle than the normal bundle of a general member $X\subset\mathbb{P}^{N}$ in the family, about which we know very little.
In this paper we study the extendability of Calabi-Yau threefolds $X\subset\mathbb{P}^{N}$ by degenerating it to Calabi-Yau ribbons on Fano-threefolds $Y$ of Picard rank one, embedded by the complete linear series of $-lH$ where $H$ is the generator of Pic$(Y)$. For $l$ equal to the index of $Y$, we show that the Calabi-Yau threefolds are at least as many times smoothly extendable as the Fano threefold $Y$. For higher values of $l$, we show non-extendability of $X$. Fano threefolds of Picard rank $1$ are classified into $17$ different deformation types by the classification of Iskovskih-Mori-Mukai ([14], [15], [34]). We use the explicit description of each of the families which are either complete intersections in weighted projective spaces or regular sections of homogeneous vector bundles on Grassmannians (see [11], [4], https://www.fanography.info/.). The computations boil down to Borel-Bott-Weil theorems. We now state our results. First we state a Proposition describing the Calabi-Yau threefolds we work with and their Hilbert schemes.
@@
    There exists a unique irreducible component $\mathcal{H}^{Y}_{l}$ of the Hilbert scheme of $\mathbb{P}^{N_{l}}$, where $N_{l}+1=h^{0}(lH)+h^{0}((l-j)H)$ with Hilbert polynomial $p(z)=h^{0}(lzH)+h^{0}((lz-j)H)$ that parameterizes linearly normal Calabi-Yau varieties $X_{t}\subset\mathbb{P}^{N_{l}}$ such that $(X_{t},\mathcal{O}_{X_{t}}(1))$ can be deformed to $(X,lA)$ along an irreducible one parameter family.
@@
It is interesting to compare the results of this article with the results on extendability of $K3$ surfaces and canonical curves in [7], [8] and [9]. Let $\mathcal{H}_{r,g}$ denote the Hilbert scheme of $K3$ surfaces embedded by the $r-$ th veronese map of a primitive very ample line bundle of sectional genus $g$. By studying Gaussian maps of canonical curve sections of such $K3$ surfaces, it was proven that for $(r=1,g\geq 13),(r=2,g\geq 7),(r=3,g\geq 5),(r=4,g\geq 4),(r\geq 5,g\geq 3)$ (see [9, Table $2.14$] for the precise and slightly stronger results), the general curve section of the general $K3$ surfaces in $\mathcal{H}_{r,g}$ has corank one Gaussian map which in particular means that the general $K3$ surfaces are not extendable and the general curve sections are exactly one extendable. The results Theorem 1.6, $(1)$, $(3)$ are analogues of the above results. However, there is a contrast for Calabi-Yau threefolds; the hyperplane sections of $\mathcal{H}^{Y}_{l}$ for $l\geq l_{Y}$ fill out an entire irreducible component of the Hilbert scheme of canonical surfaces, while curves on a $K3$ surface are non-general curves of the corresponding component of the Hilbert scheme. We plot the invariants of the surface sections in subsection 4.3.
Acknowledgements. We thank Professor Ciro Ciliberto for his encouragement.
@@ 2. Main result on extendability of projective varieties
    Ribbons $\widetilde{Y}\subset\mathbb{P}^{N}$ on $Y$ with conormal bundle $L$ extending a given (possibly degenerate) embedding $Y\subset\mathbb{P}^{N}$ are parameterized by the nowhere vanishing sections of $H^{0}(N_{Y/\mathbb{P}^{N}}\otimes L)$ (see [21, Proposition $2.1$]).
@@
Proof. We have the following exact sequences (see [17, Lemma $4.1$])
@@
Since $\alpha(X)=0$, we have that $h^{0}(N_{X/\mathbb{P}^{N+1}}(-2))=0$ (see [29, Proof of Theorem $1.3$, Claim $2.3$]). Combined with the fact that $h^{1}(N_{X/\mathbb{P}^{N+1}}(-3))=0$, we have that $h^{0}(N_{Z/\mathbb{P}^{N}})(-2)=0$. Now $h^{0}(N_{Z/\mathbb{P}^{N}})(-1)-N-1=\alpha(Z)$. Hence the dimension of the tangent space to the Hilbert scheme at the point $C(Z)$ is given by $h^{0}(N_{Z/\mathbb{P}^{N}})+\alpha(Z)+N+1=h^{0}(N_{Z/\mathbb{P}^{N}})+\alpha(X)+N+2=h^{0}(N_{Z/\mathbb{P}^{N}})+h^{0}(N_{X/\mathbb{P}^{N+1}}(-1))$.
@@ 3. Calabi-Yau ribbons on Fano threefolds and deformations of Calabi-Yau double covers of Fano threefolds
    There exists a unique irreducible component $\mathcal{H}^{Y}_{l}$ of the Hilbert scheme of $\mathbb{P}^{N_{l}}$ with Hilbert polynomial $p(z)=h^{0}(lzH)+h^{0}((lz-j)H)$ that parameterizes linearly normal Calabi-Yau varieties $X_{t}\subset\mathbb{P}^{N_{l}}$ such that $(X_{t},\mathcal{O}_{X_{t}}(1)=lA_{t})$ can be deformed to $(X,lA)$ along an irreducible one parameter family.
@@
Proof. Let us first show the values of $j$. For families $1.2-1.10$, $1.12-1.17$, this is clear by [4] since $-K_{Y}$ is very ample. For $1.1$, $H=-K_{Y}$. $Y$ is a double cover of $\mathbb{P}^{3}$ branched along a sextic. We have $h^{0}(-2K_{Y})=h^{0}(\mathcal{O}_{\mathbb{P}^{3}}(2)$ and hence the morphism induced by $|-2K_{Y}|$ is $2:1$ onto the second Veronese embedding of $\mathbb{P}^{3}$. Further one can check that $-3K_{Y}$ is in fact projectively normal. So $j=3$. For $1.11$, by [4], $j\geq 3$. Now by [35, Theorem $0.2$], $lH$ is very ample on $Y$ for $l\geq 3$, which implies $j=3$.
Now we show that if $l<j$, then $lA_{t}$ is not very ample. For family $1.1$, $l=1$, if $A_{t}$ is very ample, then it embeds $X_{t}$ as a degree $4$ Calabi-Yau threefold inside $\mathbb{P}^{4}$. This gives a contradiction. For family $1.11$, $l=1$, $A_{t}$ is not very ample since $h^{0}(A_{t})=3$. For $l=2$, if $|2A_{t}|$ is very ample, then $X_{t}$ is an aCM Calabi-Yau threefold of degree $16$ inside $\mathbb{P}^{6}$. But such a Calabi-Yau does not exist due to [24, Theorem $3.2$]. For family $1.12-1.15$, $j=2$. For $l=1$, $A_{t}$ is not very ample. To see this first note that $h^{0}(A_{t})=h^{0}(H)$. Now for $1.12$, $h^{0}(A_{t})=4$ and the morphism given by $|A_{t}|$ is $4:1$ onto $\mathbb{P}^{3}$. For $1.13-1.14$, if $A_{t}$ is very ample, then $|A_{t}|$ embeds $X_{t}$ inside $\mathbb{P}^{5}$. By [2, Remark $11$], all Calabi-Yau threefolds inside $\mathbb{P}^{5}$ are complete intersections of either two cubics or a quadric and a quartic or a quintic and a hyperplane. For $1.13$, $A_{t}^{3}=6$, which cannot be the degree of a Calabi-Yau inside $\mathbb{P}^{5}$. For $1.14$, $A_{t}^{3}=8$, so $X_{t}$ has to be the complete intersection of a quadric and a quartic. But then this is not possible by comparing $h^{1,2}(X_{t})$ with $h^{1,2}$ of the complete intersection. For $1.15$, if $A_{t}$ is very ample then $|A_{t}|$ embeds $X_{t}$ as a degree $10$, aCM Calabi-Yau threefold inside $\mathbb{P}^{6}$. By [24, Theorem $3.2$, Section $6$], such a Calabi-Yau threefold cannot exist. For family $1.16$, $j=3$. For $l=1,2$, $lA_{t}$ is not very ample. Note first that $H^{1}(T_{Y}\otimes K_{Y})=H^{1,2}(Y)=0$. Now since $h^{0}(lA)=h^{0}(lH)$, the morphism given by $|A|$ on the double cover $X$ factors through a $2:1$ map onto $Y$ followed by $Y$ embedded by $|lH|$. This together with the fact that $H^{1}(\mathcal{O}_{X})=0$, implies by [42, Corollary $1.1$] and arguments similar to [20, Theorem $1.7$] the morphism by $|lA_{t}|$ factors as a double cover of $Y$ as well. For family $1.17$, $j=4$. For $l=1,2,3$, the map given by $|A_{t}|$ is not very ample by exactly same arguments as in part $(5)$. The last statement follows by combining Lemma 3.2 and Proposition 3.1. $\blacksquare$
@@
Noting that $H^{0}(lH\otimes K_{Y})=H^{0}((l-j)H)$ we have $N^{\prime}+1=h^{0}(lH)+h^{0}((l-j)H)=N_{l}+1$. We show that all such ribbons $\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ have represent smooth points of the Hilbert scheme. This can be checked using [5, Lemma $2.14$] and the exact sequence
@@
Now we prove part $(1)$. Consider the Calabi-Yau double cover $\pi:X\to Y$ branched along $B\in|-2K_{Y}|$. By [36], Lemma $4.2$, we have $H^{1}(\Omega_{X}^{1})=H^{1}(\pi_{*}\Omega_{X}^{1})=H^{1}(\Omega_{Y})\bigoplus H^{1}(\Omega_{Y}(\log(B))\otimes K_{Y})$. Now using the exact sequence
@@
Therefore $H^{1}(\Omega_{Y}(\log(B))\otimes K_{Y})=0$ implies $\pi^{*}$ is an isomorphism, $h^{1,1}(X)=1$ and $\textrm{Pic}(X)=\mathbb{Z}$. Let $A$ denote the generator. We have that for $X_{t}$, a general deformation of $X$, $h^{1,1}(X_{t})=1$. On the other hand, by [10, Corollary 4.18] and [26, Theorem 0.1], $Y$ is simply connected, so, by the Lefschetz theorem for homotopy groups (see [28, Theorem 3.1.21]), so is $B$, where $B$ is the ample branch divisor of the double cover $X$. Now if $R\subset X$ denotes the ramification divisor, $R\cong B$ and hence $R$ is simply connected. Since $R$ is ample, once again by the Lefschetz theorem for homotopy groups, so is $X$. Therefore $X$ is simply connected and hence so is $X_{t}$. Therefore, $\textrm{Pic}(X_{t})=\mathbb{Z}$. Again since divisibility of a line bundle remains constant along a smooth family, we have that for any polarized deformation $(X_{t},A_{t})$ of $(X,A)$, $A_{t}$ generates the Picard group of $X_{t}$.
@@
As before since $Y$ is Fano and using Nakano-Vanishing theorem, one can check that both flanking terms of the previous exact sequence vanishes. This implies that $H^{1}(N_{\varphi})=0$ and deformations of $\varphi$ are unobstructed. This implies by [19, Proposition $1.4$], that there is a one parameter family $\Phi_{T}:\mathcal{X}\to\mathbb{P}^{N_{l}}_{T}$ such that $\Phi_{0}=\varphi$ while $\Phi_{t}=\varphi_{t}:\mathcal{X}_{t}=X_{t}\to\mathbb{P}^{N_{l}}$ is an embedding. Further (Im $\Phi)_{0}=\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ while for $t\neq 0$, (Im $\Phi)_{t}=X_{t}\subset\mathbb{P}^{N_{l}}$. Note that since $\widetilde{Y}\subset\mathbb{P}^{N_{l}}$ is linearly normal, we have that $X_{t}\subset\mathbb{P}^{N_{l}}$ is also linearly normal. Further we have $(\Phi_{T}^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}_{T}}(1))_{0}=\varphi^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}}(1)=\pi^{*}(lH)=lA$, and hence $(\Phi_{T}^{*}\mathcal{O}_{\mathbb{P}^{N_{l}}_{T}}(1))_{t}=\mathcal{O}_{X_{t}}(1)=lA_{t}$ is deformation of $A$.